OpenStudy (karatechopper):

@imranmeah91 help fast!!

OpenStudy (karatechopper):

\[(3^{3}\sqrt{15}\div 5\sqrt{3})^{2}\]

OpenStudy (karatechopper):

help! its in fraction form..i couldnt find a fraction bar

OpenStudy (anonymous):

3^3=27 Sqrt[15]=sqrt[5]* sqrt[3]

OpenStudy (anonymous):

well use the "/" and complete the question

OpenStudy (diyadiya):

\[\large \left( \frac{3^3\sqrt{15}}{5\sqrt{3}} \right)^2\]

OpenStudy (diyadiya):

\[\large \sqrt{15}=\sqrt{5} \times \sqrt{3}\]

OpenStudy (diyadiya):

\[\large \left( \frac{3^3\sqrt{5} {\sqrt{3}}}{5\sqrt{3}} \right)^2\]

OpenStudy (diyadiya):

\[5=\sqrt{5} \times \sqrt{5}\]

OpenStudy (diyadiya):

\[\left( \frac{3^3}{\sqrt{5}} \right)^2 = \frac{3^6}{5}\]

OpenStudy (karatechopper):

this still confuses me..

OpenStudy (diyadiya):

\[\large \left( \frac{3^3\sqrt{5} {\sqrt{3}}}{5\sqrt{3}} \right)^2\] got this much ?

OpenStudy (karatechopper):

yep

OpenStudy (diyadiya):

Sure ? then 5 can be written as \(\sqrt{5} \times \sqrt{5} \) \[\large \left( \frac{3^3\sqrt{5} {\sqrt{3}}}{\sqrt{5} \sqrt{5}\sqrt{3}} \right)^2\]

OpenStudy (karatechopper):

ok

OpenStudy (diyadiya):

Now you can cancel 1 Sqrt 5 and 1 sqrt 3 from numerator and denominator \[\large \left( \frac{3^3}{\sqrt{5}} \right)^2\]

OpenStudy (karatechopper):

did tht

OpenStudy (diyadiya):

It can be written as \[\frac{(3^3)^2}{(\sqrt{5})^2}\]

OpenStudy (karatechopper):

yes

OpenStudy (diyadiya):

Now what do you get?

OpenStudy (karatechopper):

(27)^2 / 5

OpenStudy (karatechopper):

then i get..

OpenStudy (diyadiya):

Right!

OpenStudy (diyadiya):

=D

OpenStudy (karatechopper):

729/5?

OpenStudy (diyadiya):

Yup!!!

OpenStudy (karatechopper):

:D

OpenStudy (karatechopper):

ok now i have a new question..

OpenStudy (diyadiya):

Ok.. post it as new question ..this is lagging ..

OpenStudy (karatechopper):

kk